What is the wavelength of EM radiation that ejects 2.00 -eV electrons from calcium metal, given that the binding energy is 2.71 eV? What type of EM radiation is this?

1
Expert's answer
2022-04-11T12:06:45-0400

"E-E_1=\frac{hc}{\lambda},"

"\lambda=\frac{hc}{E-E_1}=280~Å."